The electric field intensity at a point situated 4 meters from a point charge is 200N/C. If the distance is reduced to 2 meters, the field intensity will be?


1
Expert's answer
2021-06-30T16:18:35-0400
"E=k\frac{q}{r^2}\\E'=200\frac{4^2}{2^2}=800\frac{N}{C}"